OpenStudy (trojanpoem):

If f'(1) = 2 , f(1) = 4 then limit 3h/ f(1+h) -2 h - > 0

OpenStudy (freckles):

\[\lim_{h \rightarrow 0}(\frac{3h}{f(1+h)}-2) \text{ or } \lim_{h \rightarrow 0}(\frac{3h}{f(1+h)-2}) ?\] which of these is what you mean?

OpenStudy (freckles):

direct sub can be done on either one of these we won't even need f'(1)=2

OpenStudy (trojanpoem):

the second

OpenStudy (phi):

the f'(1)=2 tells us f() is differentiable at x=1, and that allows us to assume the limit f(1+h) --> f(1) as h-->0

OpenStudy (trojanpoem):

limit 3h/ f(1+h) -2 = ..... <- forgot the equal

OpenStudy (trojanpoem):

I don't know F(x) :/

OpenStudy (freckles):

have you tried that direct subsitution thingy yet?

OpenStudy (trojanpoem):

Show me what do you mean

OpenStudy (freckles):

like and yeah f'(1)=2 does actually say f is diff at x=1 which means it is continuous at x=1 so yeah phi was correct to say that is why we can assume as h->0 f(1+h)->f(1)=2

OpenStudy (freckles):

just replace the yh's with 0

OpenStudy (freckles):

h's*

OpenStudy (trojanpoem):

The book solution is 3/4

OpenStudy (freckles):

not possible

OpenStudy (trojanpoem):

Possible :D

OpenStudy (freckles):

unless you have gave the wrong problem

OpenStudy (trojanpoem):

If f was a function : F'(1) =2 , F(1) = 4 so : \[\lim_{h \rightarrow 0 } \frac{ 3h }{ F(1+h) -2 }\] = ........

OpenStudy (trojanpoem):

Solution : [3/4]

OpenStudy (phi):

oh. I guess we have to do this \[ f'(1) = \lim_{h\rightarrow 0} \frac{ f(1+h) - f(1)}{h} \]

OpenStudy (trojanpoem):

That's what I tried I got 3h/ 2 + 2h

OpenStudy (trojanpoem):

What I was suspicious of is that maybe f(1) = 2 and F'(1) =4 ? so 3h/ 4h = 3/4 but I am always wrong.

OpenStudy (phi):

yes, I think we have to assume f'(1)= 4 and f(1)=2, so we can do this \[ \lim_{h \rightarrow 0}\left( \frac{3h}{f(1+h)-2}\right) \\ = 3 \lim_{h \rightarrow 0} \frac{1}{ \frac{f(1+h)-2}{h} } \\ = 3 \lim_{h \rightarrow 0} \frac{1}{ \frac{f(1+h)-f(1)}{h} } \ \]

OpenStudy (phi):

that turns into 3 * 1/f'(1) so if f'(1) = 4 we get 3/4

OpenStudy (freckles):

I thought f'(1)=2 ?

OpenStudy (freckles):

and f(1)=4

OpenStudy (trojanpoem):

It is.

OpenStudy (freckles):

\[\lim_{h \rightarrow 0}(\frac{3h}{f(1+h)-2}) \] wait so if it was the other way around f(1)=2 and f'(1)=4 if we tried direct sub we would get 0/0 then we would apply l'hospital \[\lim_{h \rightarrow 0}\frac{3}{f'(1+h)} \\ =\frac{3}{4}\]

OpenStudy (freckles):

but the other way around it is 0

OpenStudy (phi):

yes, it looks like they switched the f and f' values. otherwise I don't see how to get the book answer.

OpenStudy (freckles):

ditto!

OpenStudy (trojanpoem):

So the book substituted them ?

OpenStudy (phi):

you can use the definition of the derivative (see above) or use L'hopital (see above)

OpenStudy (trojanpoem):

Thanks both of you guys.